OpenStudy (aaronandyson):

400/x - 400/x+12 = 5/3 Find x

OpenStudy (jhannybean):

Is your question: \(\dfrac{400}{x}-\dfrac{400}{x+12}=\dfrac{5}{3}\) ?

OpenStudy (aaronandyson):

Yes

OpenStudy (jhannybean):

Find the common denominator between x, x+12 and 3.

OpenStudy (jhannybean):

That would be \[x(x+12)(3)\]

OpenStudy (jhannybean):

So now multiply both sides of your equation by \(x(x+12)(3)\)

OpenStudy (jhannybean):

\[\large x(x+12)(3)\left[ \dfrac{400}{x}-\dfrac{400}{x+12}\right]=\left[\dfrac{5}{3}\right]x(x+12)(3)\]

OpenStudy (jhannybean):

Distribute \(x(x+12)(3)\) to each term and cancel out the like terms. This ill help you eliminate the denominator for you to solve a linear function.

OpenStudy (aaronandyson):

?

OpenStudy (jhannybean):

What is \[\frac{400}{x} \cdot x(x+12)(3)=~? \]What cancels out here?

OpenStudy (aaronandyson):

x

OpenStudy (aaronandyson):

400(x+12)(3)?

OpenStudy (jhannybean):

Great. That is correct.

OpenStudy (jhannybean):

\[-\frac{400}{x+12} \cdot x(x+12)(3)=~?\]

OpenStudy (jhannybean):

\[\frac{5}{3}\cdot x(x+12)(3)=~?\]

OpenStudy (aaronandyson):

5x(x+12)(3)?

OpenStudy (aaronandyson):

and 400x(3)

OpenStudy (jhannybean):

Awesome. Thats correct.

OpenStudy (jhannybean):

So now you have \[400(x+12)(3) -400(3)(x)=5(x)(x+12)\]

OpenStudy (jhannybean):

Do you think you can solve it from here?

OpenStudy (aaronandyson):

umm? can we cancel "x+12" from both sides?

OpenStudy (jhannybean):

No because it is not a common term on both sides of the equation. \((x+12)\) is not multiplied to \(-400(x)(3)\)

OpenStudy (jhannybean):

The only time we would be able to cancel something like that out is if we factored it out from the left side of the equation, that is, if it was a common term.

OpenStudy (aaronandyson):

What next?

OpenStudy (jhannybean):

Just expand the left hand side and the right hand side of the equation and solve for x.

OpenStudy (jhannybean):

So what do you get for the RHS of the equation first, since that's easier to solve?

OpenStudy (aaronandyson):

5x^2+60x?

OpenStudy (jhannybean):

Yup.

OpenStudy (jhannybean):

And \(400(3)(x+12)=~?\)

OpenStudy (aaronandyson):

400(3x + 36) 1200x +14400?

OpenStudy (jhannybean):

Good.

OpenStudy (jhannybean):

Putting that altogether, we get \[1200x+14400-1200x=5x^2+60x\]

OpenStudy (jhannybean):

See anything that cancels out?

OpenStudy (aaronandyson):

1200x

OpenStudy (jhannybean):

So what does our equation become when 1200x is eliminated?

OpenStudy (aaronandyson):

5x^2 + 60x - 14400 = 0

OpenStudy (jhannybean):

Good, now divide everything by 5.

OpenStudy (jhannybean):

What do you get? @AaronAndyson

OpenStudy (aaronandyson):

Sorry I had to wash my hands I got x^2 + 12 - 2880 = 0

OpenStudy (jhannybean):

Awesome, now we get to use the quadratic formula to find our roots, or x values.

OpenStudy (jhannybean):

So we solve for \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]\[x=\frac{-12\pm\sqrt{(12)^2-4(1)(-2880)}}{2}\]

OpenStudy (aaronandyson):

I'm suppose to factor -_-

OpenStudy (jhannybean):

nope. What is \((12)^2-4(-2880)=~?\)

OpenStudy (aaronandyson):

They had said that we are suppose to factor

OpenStudy (jhannybean):

Oh, you can't use the quadratic? Ok.

OpenStudy (jhannybean):

Well... What are two numbers that multiply to give you -2880 and add to give you 12?

OpenStudy (aaronandyson):

When they specific in the question to use the formula only then we can use it over here they did'nt

OpenStudy (jhannybean):

ahh. Gotcha.

OpenStudy (aaronandyson):

and 24 , 12?

OpenStudy (jhannybean):

Does 24 \(\times\) 12 = 2880?

OpenStudy (aaronandyson):

240*

OpenStudy (jhannybean):

Write out all the factors for 12, and all the factors for 2880.Maybe that will help

OpenStudy (jhannybean):

Yes, 240 \(\times \) 12 = 2880

OpenStudy (jhannybean):

But... does 240 -12 = 12?

OpenStudy (aaronandyson):

No

OpenStudy (jhannybean):

Alright, start at 40, and divide 2880 by 40. what will that give you?

OpenStudy (jhannybean):

Then take the number produced and subtract it by your factor. This will help you find the middle number

OpenStudy (aaronandyson):

72?

OpenStudy (jhannybean):

Yes, therefore does 72-40 give you 12?

OpenStudy (aaronandyson):

No

OpenStudy (jhannybean):

Move up a little bit. you know 2880 ends in a 0 therefore it has to be a number divisible by an even factor....along those lines. So try 42.

OpenStudy (aaronandyson):

68.571

OpenStudy (jhannybean):

Doesnt work. moving on. 45?

OpenStudy (aaronandyson):

64?

OpenStudy (jhannybean):

And 64 - 45 = ?

OpenStudy (aaronandyson):

19?

OpenStudy (jhannybean):

yeah, close right? Moving on.

OpenStudy (jhannybean):

Lets try another even number. 46?

OpenStudy (aaronandyson):

62.60

OpenStudy (jhannybean):

Doesnt work. Moving on once again

OpenStudy (jhannybean):

How about 48?

OpenStudy (aaronandyson):

60 and 60-48 = 12

OpenStudy (jhannybean):

Woot. there you go.

OpenStudy (jhannybean):

Now that we've found out factors, lets write our quadratic in factored form.

OpenStudy (jhannybean):

What would be the factored form with 48 and 60 as our factors?

OpenStudy (aaronandyson):

x = -48 or x = 60

OpenStudy (aaronandyson):

-60 and + 48*

OpenStudy (jhannybean):

Yep :)

OpenStudy (jhannybean):

Good job.

OpenStudy (aaronandyson):

Thanks Can you help me more?

OpenStudy (jhannybean):

Sorry, its kind of late here and I'm about to head off.

OpenStudy (aaronandyson):

Then,Goodnight.

OpenStudy (jhannybean):

Good luck

OpenStudy (aaronandyson):

Thanks